OpenStudy (anonymous):

DUE TOMORROW, THIS PROBLEM HAS BEEN GIVING ME TROUBLE ALL DAY! •• Really stuck with this problem, not sure what to do: Consider a system where the three fundamentally important quantities are the speed of light c with dimensions [L]/[T], Planck's constant h with dimensions [M][L]^2/[T] and the mass of the proton mp with dimension [M]. What combination of ratios and/or products of c, h, and mp will yield a new quantity of dimensions [T]?

OpenStudy (anonymous):

If you can help, please help me! I posted this question on OpenStudy for 6 hours!

OpenStudy (anonymous):

\[h \times \frac{ 1 }{ mp }\times \frac{ 1 }{ c }\]

OpenStudy (anonymous):

? I'm stuck on even how you get started with these problems... What's the question the problem is trying to ask?

OpenStudy (anonymous):

@Euler271

OpenStudy (anonymous):

Sort of like this: http://openstudy.com/updates/5254715ae4b07dfb515fe9de except Im stuck for [T]